OpenStudy (anonymous):

solve the following equation 5(z-4)-z=4z-20

OpenStudy (jhannybean):

expand the LHS, what do you get?

OpenStudy (anonymous):

what does LHS stand for?

OpenStudy (sleepyhead314):

Hello and Welcome to OpenStudy! :) to solve for z first distribute the 5 can you do that? what would you get?

OpenStudy (anonymous):

5z-20?

OpenStudy (sleepyhead314):

yes :) good! then add like terms 5z - 20 - z = 4z - 20

OpenStudy (sleepyhead314):

then... oh I see why you might be confused... ok then if you get 0=0 (a true statement) then there are in infinite number of solutions if you get something like 0=-20 (a false statement) then there are no solutions

OpenStudy (jhannybean):

LHS stands for Left Hand Side. Have you added the like terms?

OpenStudy (skullpatrol):

The equation has infinitely many solutions. You can write this as the solution set is the set of all real numbers. Because any number you substitute in for z will give you a true statement and thus satisfy the equation and be apart of its solution set, just like the equation y+2=2+y is true for all real numbers.